OpenStudy (anonymous):

The functions u and w are defined as follows. u(x)=-x-2 w(x)=-2x^(exponent 2)+1 Find the value of u(w(4))

OpenStudy (anonymous):

No clue :/

OpenStudy (anonymous):

Okay Not here to answer your Q just here to say

OpenStudy (anonymous):

Yes i think i know how to solve a function

OpenStudy (anonymous):

You cant really tell me what to do... All you had to do ws let me finnish you not my gaurdain

OpenStudy (solomonzelman):

\(\large\color{red}{ \displaystyle u(x)=-x-2 }\) \(\large\color{blue}{ \displaystyle w(x)=-2x^2+1 }\) Find the value of u(w(4)) first, can you find \(\large\color{black}{ \displaystyle \color{red}{u}(~\color{blue}{w(x)}~) }\) ?

OpenStudy (solomonzelman):

(plug in what w(x) is equal to, into u(x) instead of x.)

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle \color{red}{u}(~\color{blue}{w(x)}~) = \color{red}{-\color{blue}{(-2x^2+1)}-2}}\)

OpenStudy (solomonzelman):

then as you simplify the \(\large\color{black}{ \displaystyle \color{red}{u}(~\color{blue}{w(x)}~) }\) and plug in 4. \(\large\color{black}{ \displaystyle (}\)or ever, right away plug in and then simplify\(\large\color{black}{ \displaystyle )}\). this will give you \(\large\color{black}{ \displaystyle \color{red}{u}(~\color{blue}{w(4)}~)}\)